OpenStudy (anonymous):

Please help! I'll give you a medal. Simplify the expression 4x^5y^-4/ -2x^3y^4

OpenStudy (anonymous):

when dividing monomials, Subtract the exponents For example\[\frac{ x^5 }{ x^3 }=x^{5-3}=x^2\] What do you get when you divide the y's?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

I just need the answer. On a test.